1. Given the demand function as;

and the supply function as;

Also the Marginal cost is given by and the fixed cost is 55
a) If the firm decides to maximise total revenue, fine the consumer’ surplus
b) If the firm decides to maximise profit, fine the consumer’ surplus
c) Find the producers’ surplus at the market equilibrium
1
Expert's answer
2020-06-01T12:58:45-0400

a) If the firm decides to maximise total revenue, then P = Q, and the consumer’ surplus is "CS = 0.5*(P0 - P)*Q" , where P0 is price at Q = 0.

b) If the firm decides to maximise profit, the firm will produce the quantity for which MR = MC at price from the demand curve, and the consumer’ surplus is "CS = 0.5*(P0 - P)*Q" , where P0 is price at Q = 0.

c) The market equilibrium is at MC = D, the producers’ surplus is PS = 0.5*(P - MC0)*Q, where MC0 is MC at Q = 0.